# Topology optimization of nonlinear forced response curves via reduction on spectral submanifolds

### v1

@@ 1 Introduction
Nonlinear dynamic responses are commonly observed in micro-electro-mechanical systems (MEMS). On the one hand, such responses may cause the systems to fail to work properly. For example, in some high-precision sensors, such as gyroscopes su_characteristics_2020 ; tiwari_using_2019 , pressure sensors LI201219 , nonlinear responses reduce the accuracy of the sensor and even cause malfunctions s20247207 . On the other hand, nonlinearities can also be utilized to provide advantages over linear designs. For example, nonlinear dynamic responses can be used to improve the performance of energy harvesters chen2015internal ; zhang2024internal ; amin_karami_powering_2012 , and their characteristic hardening or softening behavior can be utilized in the design of filters hajjaj_mode_2017 and binary counters yao_counter_2013 . Thus, it is essential to tailor the dynamic responses of nonlinear systems if one aims to suppress the amplitude of structural vibration or to utilize them to optimally design devices.
@@
To utilize nonlinear dynamic responses, several studies have focused on tailoring the hardening and softening behavior of nonlinear systems. One approach involves adjusting the hardening behavior of clamped-clamped beams by maximizing or minimizing the coefficient of cubic nonlinearity in the normal form dou_structural_2015 . Another method optimizes both the Q-factor and Duffing nonlinearity in finite element-based reduced-order models using derivative-free optimization techniques li_finite_2025 . Additionally, backbone curves have been optimized on a reduced-order model (ROM) constructed via Lyapunov subcenter manifolds pozzi_topology_2025 ; pozzi_backbone_2024 and spectral submanifolds (SSMs) pozzi2025adjointsensitivitiesoptimizationnonlinear .
@@
Tailoring the hardening and softening behavior of a nonlinear system typically involves analyzing its autonomous dynamics and does not require consideration of external forcing. However, if the objective is to suppress or enhance the vibration amplitude under harmonic excitation, it is necessary to tailor the forced response curve (FRC). For low-dimensional nonlinear systems, the FRC of nonlinear systems can be computed by the harmonic balance method von_groll_harmonic_2001 ; krack_application_2019 ; detroux_harmonic_2015 combined with continuation techniques ascher1995numerical ; dankowicz2013recipes . Based on these techniques, dou_optimization_2015 applied shape optimization to a clamped-clamped beam to minimize or maximize the amplitude of primary resonance. For high-dimensional nonlinear systems, such as finite element models (FEM) in topology optimization bendsoe2013topology , computing and tailoring the FRC becomes challenging due to the increased computational cost.
@@
In addition to amplitude control, tailoring the bifurcations of the FRC is also essential to nonlinear systems. Under harmonic excitation, nonlinear systems often exhibit saddle-node (SN) bifurcations on the FRC as the forcing amplitude increases. These bifurcations lead to abrupt jumps in the steady-state response, known as jump or hysteresis phenomena, which are particularly detrimental to devices intended to operate within a linear or predictable regime. Therefore, suppressing the occurrence of SN bifurcations becomes critical. In previous work, shape optimization along with numerical continuation is used to tailor the distance of two SN bifurcations on the nonlinear FRC saghafi_nonlinear_2015 .
@@
Here, we aim to use topology optimization to tune the FRC of high-dimensional nonlinear systems, including minimizing or maximizing the peak of FRC, manipulating the hardening/softening behavior, and suppressing the occurrence of SN bifurcations. As a method to design continuum structures noauthor_topology_2004 , topology optimization can produce more refined structures than shape optimization. However, the number of degrees of freedom (DOFs) of the FEM and the associated design variables in topology optimization is considerable.
@@
To address the challenge of high-dimensionality, reduction on SSMs has been used to analytically extract the FRC of FEM with high DOFs haller_nonlinear_2016 ; li_nonlinear_2022 ; haller_modeling_2025 ; kaszas_globalizing_2025 . In particular, SSM reduction transforms forced periodic orbits into fixed points of the associated reduced-order model (ROMs). In addition, the sensitivity of coefficients in SSM-based ROMs that govern the peak of FRC and hardening/softening behavior has been discussed in li_explicit_2024 . These sensitivities and their efficient computation pozzi_topology_2025 are essential in the iteration process of topology optimization. Based on these contributions, we will perform topology optimization of FRCs.
@@
In this work, we adopt a density-based topology optimization algorithm to determine the optimal material layout. Specifically, the Solid Isotropic Material with Penalization (SIMP) scheme bendsoe_material_1999 is used to interpolate material properties and promote 0–1 designs. To enhance numerical stability and obtain clear structural boundaries, both density filtering and density projection techniques are employed wang_projection_2011 . The design variables are updated using the Method of Moving Asymptotes (MMA) svanberg_method_1987 . All these techniques follow the implementation described in pozzi_topology_2025 .
@@ 2.2 Reduction on spectral submanifolds
In this study, we construct the two-dimensional SSM associated with the master subspace $\mathcal{E}=\mathrm{span}(\boldsymbol{v}_{1},\bar{\boldsymbol{v}}_{1})$ for model reduction. Specifically, let $\boldsymbol{p}=(p,\bar{p})$ denote the reduced coordinates. We seek the autonomous SSM map $\boldsymbol{z}=\boldsymbol{W}(\boldsymbol{p})$ and its associated reduced dynamics $\dot{\boldsymbol{p}}=\boldsymbol{R}(\boldsymbol{p})$ at $\epsilon=0$. When external periodic forcing $\epsilon\boldsymbol{f}^{\mathrm{ext}}(\Omega t)$ is considered, the SSM becomes time-dependent, yielding a periodic SSM map $\boldsymbol{W}_{\epsilon}(\boldsymbol{p},\Omega t)$ and reduced dynamics $\dot{\boldsymbol{p}}=\boldsymbol{R}_{\epsilon}(\boldsymbol{p},\Omega t)$. The map and reduced dynamics on the periodic SSM can be expressed as li_nonlinear_2022 ; li_explicit_2024
@@
Cohomological equations can be derived, and the expansion coefficients can be obtained following a normal-form-style parameterization. More details can be found at li_explicit_2024 .
@@
One can again use the invariance equation to solve for these unknown coefficients. Interested readers can refer to li_explicit_2024 for more details.
@@
and plug it into the reduced dynamics in (9). Then the third-order reduced dynamics can be expressed as li_explicit_2024
@@
where $\lambda$ is the eigenvalue of master subspace, i.e., $\lambda_{1}$; $\gamma$ is the third-order backbone coefficient; $\tilde{f}$ is a modal force. Explicit expressions for $\gamma$ and $\tilde{f}$ are given as
@@
where $\boldsymbol{f}_{21}$ is associated with cubic nonlinearity, given in the Eq.(19) of li_explicit_2024 . It is noted that a fixed point of the ROM (15) corresponds to a forced periodic orbit of the full system (1), and the forced response curve (FRC) of the full system can be analytically extracted via the SSM-based ROM li_nonlinear_2022 ; li_explicit_2024 .
@@ 2.3 Optimization problem statement
In this paper, we consider a few optimization problems associated to the tuning of the FRC. As we will see, these optimization problems enable the tuning of the peak and the hardening/softening behavior of FRC, and also the SN bifurcations on the FRC.
@@ 2.3.1 Minimize the peak of FRC with target backbone
Since the fixed point of the reduced dynamics (15) corresponds to a periodic orbit of the full system li_nonlinear_2022 ; li_explicit_2024 , setting $(\dot{\rho},\dot{\theta})=\boldsymbol{0}$ and eliminating $\theta$, we obtain li_explicit_2024
@@
This algebraic equation gives the FRC in reduced coordinates $\rho(\Omega)$, which can be further mapped to the FRC of periodic orbits of the full system via the periodic SSM map $\boldsymbol{W}_{\epsilon}(\boldsymbol{p},\Omega t)$ in (9).
Furthermore, the peak point $(\rho_{\rm{max}},\Omega_{\rm{max}})$ on the reduced FRC $\rho(\Omega)$ corresponds to the peak of FRC in physical coordinates li_explicit_2024 ; breunung_explicit_2018 . Let $d\rho/d\Omega=0$, the peak point in reduced coordinates can be expressed as li_explicit_2024
@@
Equation (18) provides the analytical expression of the backbone curve in reduced coordinates. Specifically, when $\mathrm{Im}(\gamma)>0$, we obtain $\Omega_{\rm{max}}>\mathrm{Im}(\lambda)$, corresponding to a hardening behavior; when $\mathrm{Im}(\gamma)<0$, we obtain $\Omega_{\rm{max}}<\mathrm{Im}(\lambda)$, corresponding to a softening behavior. The corresponding amplitude $\rho_{\rm{max}}$ can be computed by solving Equation (19).
@@
Here, the constraint on $\mathrm{Im}(\gamma)$ is added to control the hardening/softening behavior. We also add constraints on $\omega_{\text{Y}}$ (i.e., $\omega_{1}$) and $\omega_{\text{X}}$ (i.e., $\omega_{2}$) to prevent the occurrence of internal resonance in the first and second modes. Specifically, the condition $\omega_{X,\text{target}}>3\omega_{Y,\text{target}}$ needs to be satisfied since the highest-order nonlinear term in the system is cubic. Further, the structural area fraction $A$ is subject to an upper bound.
@@
For the first reference problem, we consider a linear optimization formulation for periodic response dou_optimization_2015 ; niu_objective_2018 , namely
@@
where $\boldsymbol{U}$ denotes the linear displacement vector obtained from $(\boldsymbol{K}+\mathrm{i}\Omega\boldsymbol{C}-\Omega^{2}\boldsymbol{M})\boldsymbol{U}=\boldsymbol{f}^{\text{ext}}$, and the symbol $|\cdot|$ represents taking the absolute value of each component of the vector. Here, $\boldsymbol{L}$ is a diagonal selection matrix with ones at DOFs corresponding to the node, line, or domain of interest niu_objective_2018 . To maintain consistency, (21) has the same constraints as (20), except for the constraint on $\mathrm{Im}(\gamma)$, since linear optimization cannot capture or control the hardening/softening behavior. We will give an example to compare the optimization result of formulations (20) and (21) in Sec. 5.1.
@@
It is clear from the objective function that $c_{\gamma}=0$ is achieved when $\mathrm{Im}(\gamma)=\gamma_{\text{target}}$. The constraints in the formulation (22) are consistent with those in (20), ensuring a fair comparison between the two formulations. We will give an example to compare the result of formulations 20 and 22 in Sec. 5.2.
@@
We consider the reference problem (22) to tune the backbone curve. Since forcing is not involved in the problem, it has no effect on the peak of FRC. The regulation of softening/hardening behavior through the manipulation of backbone curve has been demonstrated in pozzi_topology_2025 ; pozzi2025adjointsensitivitiesoptimizationnonlinear . The formulation (22) is different from that of pozzi_topology_2025 in two perspectives: 1) we manipulate damped backbone curves breunung_explicit_2018 instead of undamped backbone curves as in pozzi_topology_2025 ; and 2) we set target value of $\gamma$ in the objective function instead of a constraint as in pozzi_topology_2025 . In addition, our formulation (22) tunes the backbone curve in reduced coordinates while the backbone curve is tailored by defining target points in terms of response frequency and physical amplitude in pozzi2025adjointsensitivitiesoptimizationnonlinear .

### v2

@@ 1 Introduction
Nonlinear dynamic responses are commonly observed in micro-electro-mechanical systems (MEMS). On the one hand, such responses may cause the systems to fail to work properly. For example, in some high-precision sensors, such as gyroscopes Su et al. (2020); Tiwari and Candler (2019), pressure sensors Li et al. (2012), nonlinear responses reduce the accuracy of the sensor and even cause malfunctions Li et al. (2020). On the other hand, nonlinearities can also be utilized to provide advantages over linear designs. For example, nonlinear dynamic responses can be used to improve the performance of energy harvesters Chen and Jiang (2015); Zhang et al. (2024); Amin Karami and Inman (2012), and their characteristic hardening or softening behavior can be utilized in the design of filters Hajjaj et al. (2017) and binary counters Yao and Hikihara (2013). Thus, it is essential to tailor the dynamic responses of nonlinear systems if one aims to suppress the amplitude of structural vibration or to utilize them to optimally design devices.
@@
To utilize nonlinear dynamic responses, several studies have focused on tailoring the hardening and softening behavior of nonlinear systems. One approach involves adjusting the hardening behavior of clamped-clamped beams by maximizing or minimizing the coefficient of cubic nonlinearity in the normal form Dou et al. (2015). Another method optimizes both the Q-factor and Duffing nonlinearity in finite element-based reduced-order models using derivative-free optimization techniques Li et al. (2025). Additionally, backbone curves have been optimized on a reduced-order model (ROM) constructed via Lyapunov subcenter manifolds Pozzi et al. (2025b, 2024) and spectral submanifolds (SSMs) Pozzi et al. (2025a).
@@
Tailoring the hardening and softening behavior of a nonlinear system typically involves analyzing its autonomous dynamics and does not require consideration of external forcing. However, if the objective is to suppress or enhance the vibration amplitude under harmonic excitation, it is necessary to tailor the forced response curve (FRC). For low-dimensional nonlinear systems, the FRC of nonlinear systems can be computed by the harmonic balance method Von Groll and Ewins (2001); Krack and Gross (2019); Detroux et al. (2015) combined with continuation techniques Ascher et al. (1995); Dankowicz and Schilder (2013). Based on these techniques, Dou and Jensen (2015) applied shape optimization to a clamped-clamped beam to minimize or maximize the amplitude of primary resonance. For high-dimensional nonlinear systems, such as finite element models (FEM) in topology optimization Bendsoe and Sigmund (2013), computing and tailoring the FRC becomes challenging due to the increased computational cost. In previous work, the moving morphable components (MMC) optimization framework has been employed to reduce the level of nonlinear vibrations in under-platform dampers Denimal et al. (2021).
@@
In addition to amplitude control, tailoring the bifurcations of the FRC is also essential to nonlinear systems. Under harmonic excitation, nonlinear systems often exhibit saddle-node (SN) bifurcations on the FRC as the forcing amplitude increases. These bifurcations lead to abrupt jumps in the steady-state response, known as jump or hysteresis phenomena, which are particularly detrimental to devices intended to operate within a linear or predictable regime. Therefore, suppressing the occurrence of SN bifurcations becomes critical. In previous work, shape optimization along with numerical continuation is used to tailor the distance of two SN bifurcations on the nonlinear FRC Saghafi et al. (2015). Multi-parametric continuation methods have also been used to track SN bifurcations with respect to multiple system parameters, enabling the identification of parameter ranges in which isolated solutions are suppressed Grenat et al. (2019). In addition, multi-parametric optimization frameworks have been proposed to control bifurcation behavior by regulating both the existence and the locations of different bifurcation types, including SN, branch-point, Neimark–Sacker, and period-doubling bifurcations Mélot et al. (2024a, b).
Internal resonance is another characteristic phenomenon of nonlinear systems and can lead to complex dynamic behaviors Liang et al. (2025). Topology optimization has been employed to tailor the eigenfrequencies of structures in order to avoid internal resonances Pedersen (2005); Sun (2020). In addition, shape optimization has been used to regulate energy transfer between modes involved in internal resonance, such as the 1:2 resonance case Dou et al. (2015). For systems exhibiting 1:3 internal resonance, a synthesis approach has been proposed to suppress modal interactions by translating the associated nonlinear normal mode branch toward lower energy levels Detroux et al. (2021).
@@
Here, we aim to use topology optimization to tune the FRC of high-dimensional nonlinear systems, including minimizing or maximizing the peak of FRC, manipulating the hardening/softening behavior, and suppressing the occurrence of SN bifurcations. In this study, internal resonance effects are not considered. Extending the proposed topology optimization framework to account for internal resonance phenomena constitutes an interesting direction for future research. As a method to design continuum structures Jung and Gea (2004), topology optimization can produce more refined structures than shape optimization. However, the number of degrees of freedom (DOFs) of the FEM and the associated design variables in topology optimization is considerable.
@@
To address the challenge of high-dimensionality, reduction on SSMs has been used to analytically extract the FRC of FEM with high DOFs Haller and Ponsioen (2016); Jain and Haller (2022); Haller (2025); Kaszás and Haller (2025). In particular, SSM reduction transforms forced periodic orbits into fixed points of the associated reduced-order model (ROMs). Consequently, one can obtain orders-of-magnitude speed-up gain from SSM reduction in comparison with collocation, harmonic balance, and shooting methods applied to full model Jain and Haller (2022); Li et al. (2022). In addition, the sensitivity of coefficients in SSM-based ROMs that govern the peak of FRC and hardening/softening behavior has been discussed in Li (2024). These sensitivities and their efficient computation Pozzi et al. (2025b) are essential in the iteration process of topology optimization. Based on these contributions, we will perform topology optimization of FRCs.
@@
In this work, we adopt a density-based topology optimization algorithm to determine the optimal material layout. Specifically, the Solid Isotropic Material with Penalization (SIMP) scheme Bendsøe and Sigmund (1999) is used to interpolate material properties and promote 0–1 designs. To enhance numerical stability and obtain clear structural boundaries, both density filtering and density projection techniques are employed Wang et al. (2011). The design variables are updated using the Method of Moving Asymptotes (MMA) Svanberg (1987). All these techniques follow the implementation described in Pozzi et al. (2025b).
@@ 2.1 Setup
Substitution of $\boldsymbol{\phi}_{1}^{\mathrm{T}}\boldsymbol{C}\boldsymbol{\phi}_{1}=2\xi\omega_{1}$ and Eq. (3) into Eq. (2) yields
$\lambda_{1}^{2}+2\xi\omega_{1}\lambda_{1}+\omega_{1}^{2}=0,\implies\lambda_{1}=-\xi\,\omega_{1}+\mathrm{i}\,\sqrt{1-\xi^{2}}\omega_{1}.$ (5)
@@ 2.2 Reduction on spectral submanifolds
In this study, we construct the two-dimensional SSM associated with the master subspace $\mathcal{E}=\mathrm{span}(\boldsymbol{v}_{1},\bar{\boldsymbol{v}}_{1})$ for model reduction. Specifically, let $\boldsymbol{p}=(p,\bar{p})$ denote the reduced coordinates. We seek the autonomous SSM map $\boldsymbol{z}=\boldsymbol{W}(\boldsymbol{p})$ and its associated reduced dynamics $\dot{\boldsymbol{p}}=\boldsymbol{R}(\boldsymbol{p})$ at $\epsilon=0$. When external periodic forcing $\epsilon\boldsymbol{f}^{\mathrm{ext}}(\Omega t)$ is considered, the SSM becomes time-dependent, yielding a periodic SSM map $\boldsymbol{W}_{\epsilon}(\boldsymbol{p},\Omega t)$ and reduced dynamics $\dot{\boldsymbol{p}}=\boldsymbol{R}_{\epsilon}(\boldsymbol{p},\Omega t)$. The map and reduced dynamics on the periodic SSM can be expressed as Li et al. (2022); Li (2024)
@@
Cohomological equations can be derived, and the expansion coefficients can be obtained following a normal-form-style parameterization. More details can be found at Li (2024).
@@
One can again use the invariance equation to solve for these unknown coefficients. Interested readers can refer to Li (2024) for more details.
@@
and plug it into the reduced dynamics in (10). Then the third-order reduced dynamics can be expressed as Li (2024)
@@
where $\lambda$ is the eigenvalue of master subspace, i.e., $\lambda=\lambda_{1}={\color[rgb]{0,0,1}-\xi\,\omega_{1}+i\,\sqrt{1-\xi^{2}}\omega_{1}}$, as seen in (5); $\gamma$ is the third-order backbone coefficient; $\tilde{f}$ is a modal force. Explicit expressions for $\gamma$ and $\tilde{f}$ are given as
@@
where $\boldsymbol{f}_{21}$ is associated with cubic nonlinearity, given in the Eq.(19) of Li (2024). It is noted that a fixed point of the ROM (16) corresponds to a forced periodic orbit of the full system (1), and the forced response curve (FRC) of the full system can be analytically extracted via the SSM-based ROM Li et al. (2022); Li (2024).
Remark 1
The reduced-order model in Eq. (16) is based on a third-order SSM truncation and is therefore valid only within a range of forcing amplitudes for which the third-order truncation already converges. In practice, one should check the convergence of the optimal FRC with increasing truncation orders to ensure the reliability of the SSM prediction. We will apply this checking procedure in all examples.
@@ 2.3 Optimization problem statement
In this paper, we consider a few optimization problems associated to the tuning of the FRC. As we will see, these optimization problems enable the tuning of the peak and the hardening/softening behavior of FRC, and also the SN bifurcations on the FRC (cf. Fig. 1).
Figure 1: FRC and backbone curve of the SSM-based ROM in Eq. (18). The blue solid line represents the stable segment of the FRC, the blue dashed line represents the unstable segment of the FRC, and the circle on the FRC marks the saddle-node (SN) bifurcation. The associated backbone curve is represented by the red dashed line.
@@ 2.3.1 Minimize the peak of FRC with target backbone
Since the fixed point of the reduced dynamics (16) corresponds to a periodic orbit of the full system Li et al. (2022); Li (2024), setting $(\dot{\rho},\dot{\theta})=\boldsymbol{0}$ and eliminating $\theta$, we obtain Li (2024)
@@
This algebraic equation gives the FRC in reduced coordinates $\rho(\Omega)$ as shown in Fig. 1. This FRC can be further mapped to the FRC of periodic orbits of the full system via the periodic SSM map $\boldsymbol{W}_{\epsilon}(\boldsymbol{p},\Omega t)$ in (10).
Furthermore, the peak point $(\rho_{\rm{max}},\Omega_{\rm{max}})$ on the reduced FRC $\rho(\Omega)$ corresponds to the peak of FRC in physical coordinates Li (2024); Breunung and Haller (2018). Let $d\rho/d\Omega=0$, the peak point in reduced coordinates can be expressed as Li (2024)
@@
Equation (19) provides the analytical expression of the backbone curve in reduced coordinates. Specifically, when $\mathrm{Im}(\gamma)>0$, we obtain $\Omega_{\rm{max}}>\mathrm{Im}(\lambda)$, corresponding to a hardening behavior as shown in Fig. 1; when $\mathrm{Im}(\gamma)<0$, we obtain $\Omega_{\rm{max}}<\mathrm{Im}(\lambda)$, corresponding to a softening behavior. The corresponding amplitude $\rho_{\rm{max}}$ can be computed by solving Equation (20).
@@
Here, the constraint on $\mathrm{Im}(\gamma)$ is imposed to control the hardening/softening behavior. The constraint on $\omega_{Y}$ (i.e., the first natural frequency $\omega_{1}$ associated to the vibration mode along vertical direction) is introduced to maintain sufficient stiffness in the $Y$ direction throughout the optimization process, yielding well-connected optimal topology layout (cf. Remark 4), while the constraint on $\omega_{X}$ (i.e., the second natural frequency $\omega_{2}$ associated to the vibration mode along horizontal direction) is imposed to prevent the occurrence of internal resonance between the first and second modes. Specifically, the condition $\omega_{X,\text{target}}>3\omega_{Y,\text{target}}$ must be satisfied, since the highest-order nonlinear term in the system is cubic. Further, the structural area fraction $A$ is subject to an upper bound.
@@
For the first reference problem, we consider a linear optimization formulation for periodic response Dou and Jensen (2015); Niu et al. (2018), namely
@@
where $\boldsymbol{U}$ denotes the linear displacement vector obtained from $(\boldsymbol{K}+\mathrm{i}\Omega\boldsymbol{C}-\Omega^{2}\boldsymbol{M})\boldsymbol{U}=\boldsymbol{f}^{\text{ext}}$, and the symbol $|\cdot|$ represents taking the absolute value of each component of the vector. Here, $\boldsymbol{L}$ is a diagonal selection matrix with ones at DOFs corresponding to the node, line, or domain of interest Niu et al. (2018). To maintain consistency, (22) has the same constraints as (21), except for the constraint on $\mathrm{Im}(\gamma)$, since linear optimization cannot capture or control the hardening/softening behavior. We will give an example to compare the optimization result of formulations (21) and (22) in Sec. 5.1.
@@
It is clear from the objective function that $c_{\gamma}=0$ is achieved when $\mathrm{Im}(\gamma)=\gamma_{\text{target}}$. The constraints in the formulation (23) are consistent with those in (21), ensuring a fair comparison between the two formulations. We will give an example to compare the result of formulations 21 and 23 in Sec. 5.2.
@@
We consider the reference problem (23) to tune the backbone curve. Since forcing is not involved in the problem, it has no effect on the peak of FRC. The regulation of softening/hardening behavior through the manipulation of backbone curve has been demonstrated in Pozzi et al. (2025b, a). The formulation (23) is different from that of Pozzi et al. (2025b) in two perspectives: 1) we manipulate damped backbone curves Breunung and Haller (2018) instead of undamped backbone curves as in Pozzi et al. (2025b); and 2) we set target value of $\gamma$ in the objective function instead of a constraint as in Pozzi et al. (2025b). In addition, our formulation (23) tunes the backbone curve in reduced coordinates while the backbone curve is tailored by defining target points in terms of response frequency and physical amplitude in Pozzi et al. (2025a).